OpenStudy (lifeisadangerousgame):

Help with slope again? I need to figure out how to graph a line if I only have something like this y =4x+2 for example

OpenStudy (lifeisadangerousgame):

I don't understand...

OpenStudy (johnweldon1993):

y = mx + b m = slope b is y-intercept

OpenStudy (johnweldon1993):

the y-intercept is where you put your first point....in your case at y = 2 you would put a point okay?

OpenStudy (lifeisadangerousgame):

Okay

OpenStudy (johnweldon1993):

maybe a side by side comparison will help too y =4x+2 y =mx + b so YOUR slope is 4...and y-intercept is 2 so after you put your point at y = 2 you deal with the slope

OpenStudy (johnweldon1993):

would you agree that .....4...can be written as 4/1?

OpenStudy (lifeisadangerousgame):

Yes

OpenStudy (johnweldon1993):

okay good.....now remember slope = rise / run rise (where you go in the y-direction) run (where you go in the x-direction) you have slope as 4/1 right? so if slope is rise 4 ---- and you have ---- run 1 how far in the y-direction will you travel?...and then how far in the x-direction?

OpenStudy (lifeisadangerousgame):

you travel 1 in the x-direction and 4 in the y-direction?

OpenStudy (johnweldon1993):

correct! so now starting at the point you plotted...at y = 2 you go up 4 units.....then over to the right 1 unit..... ****4 up because it is a positive 4 right? and over to the right 1 because it is also positive right?**** then you can make a point where that ends...and then do it again....and eventually...connect those dots and make your line

OpenStudy (lifeisadangerousgame):

Okay is this right? It doesn't seem like it >.<

OpenStudy (lifeisadangerousgame):

@johnweldon1993

OpenStudy (johnweldon1993):

yeah that is correct...why doesn't it seem right?

OpenStudy (lifeisadangerousgame):

I think it's my line :b It's so crooked >.<

OpenStudy (johnweldon1993):

 lol better?

OpenStudy (johnweldon1993):

actually no yours is wrong...you only went up 3 in the y-direction

OpenStudy (lifeisadangerousgame):

I should'vegone up how many instead?

OpenStudy (johnweldon1993):

but you had the right idea....remember you were supposed to go up 4?

OpenStudy (lifeisadangerousgame):

Ohh that's right

OpenStudy (johnweldon1993):

" starting at the point you plotted...at y = 2 you go up 4 units.....then over to the right 1 unit..... ****4 up because it is a positive 4 right? and over to the right 1 because it is also positive right?**** " remember that?

OpenStudy (lifeisadangerousgame):

Yes, I don't know why I didn't go four :/

OpenStudy (lifeisadangerousgame):

Do you think you could help me with another one?

OpenStudy (johnweldon1993):

no problem....but you do have the general idea!

OpenStudy (lifeisadangerousgame):

The equation is y=2x+3 this time and this is how I graphed it...also I was wondering, what do I do to complete the rest of the graph? like do I do it negatively? so go down 3 and go 2 to the left?

OpenStudy (johnweldon1993):

take it step by step......what are you supposed to do first? to determine where you start on the graph?

OpenStudy (lifeisadangerousgame):

Yes, so I determined that I start at 3 since that's the y intercept ooh :b I need to start at 3 first

OpenStudy (johnweldon1993):

haha there you go lol...and after that you dont worry about that 3 anymore do you? you ONLY care about the slope...which is?

OpenStudy (lifeisadangerousgame):

Which is 2/1 so now I got up 2 and go to the right 1 :D

OpenStudy (johnweldon1993):

correct....... now...before you had the right idea...and good job recognizing that... if you wanted to go in reverse the slope being 2/1 what is -2/-1? the SAME thing right? just in a different direction so yes you can go down 2 and to the left 1 as well to make more points

OpenStudy (lifeisadangerousgame):

So it's this :D

OpenStudy (johnweldon1993):

perfect :)

OpenStudy (lifeisadangerousgame):

Okay can we do 1 last one?

OpenStudy (johnweldon1993):

sure why not lol...but you're getting good at it

OpenStudy (lifeisadangerousgame):

:D Thanks :D This time with a slope of y=1/5x+4

OpenStudy (lifeisadangerousgame):

I know first I go up 4

OpenStudy (lifeisadangerousgame):

Wait this is just like all the other ones lol

OpenStudy (johnweldon1993):

you mean you plot a point at 4 right?

OpenStudy (lifeisadangerousgame):

Only I start with a fraction, so I go up 1 and move to the right 5 right? And yes, I mean put a point at 4 but I have to go up 4 to do that so I said go up 4 :b I should've said plot a point at 4 though xD

OpenStudy (johnweldon1993):

and yes it's just like the other ones and yes you are correct about the slope as well....so the graph will look like?

OpenStudy (lifeisadangerousgame):

Like this :D

OpenStudy (johnweldon1993):

you are now a master :D

OpenStudy (lifeisadangerousgame):

Whoop! :D

OpenStudy (lifeisadangerousgame):

I know I keep saying this..could we do one last one? I promise it's the last one :b

OpenStudy (johnweldon1993):

lol great job and feel free to message me or fan me if you need anymore help :)

OpenStudy (johnweldon1993):

or just ask another one :P lol sure :)

OpenStudy (lifeisadangerousgame):

:D Thanks haha, this time it's 3x+4y=8 so I'm not sure how to graph it

OpenStudy (johnweldon1993):

ha....this one...you need to work at it first...this is what we call the "standard form" of a line you need to convert it into slope intercept form so change 3x+4y=8 to y = mx + b can you?

OpenStudy (johnweldon1993):

****hint...notice how y is by itself in the equation you want****

OpenStudy (lifeisadangerousgame):

So...it'll be 3x + 4y = 8 -3x -3x 4y = -3x + 8 /4 /4 y = -3x/4 + 8/4 y = -3/4x + 2 is that right?

OpenStudy (johnweldon1993):

again...you are a master :)

OpenStudy (lifeisadangerousgame):

:D Math isn't so bad..yet :b

OpenStudy (johnweldon1993):

lol no of course it isn't :D

OpenStudy (lifeisadangerousgame):

since 3 is negative, do I go down 3 instead?

OpenStudy (johnweldon1993):

here's a hint try graphing it with the -3 and the / 4 and then try it with +3 and the / -4 what do you get?

OpenStudy (lifeisadangerousgame):

This is with -3 and the /4

OpenStudy (johnweldon1993):

.......waiting for it .....

OpenStudy (johnweldon1993):

lol...*****suddenly a realization will come over you*****........lol